# TARS [Uncaptioned image]: MinMax Token-Adaptive Preference Strategy for MLLM Hallucination Reduction

### v3

@@ 3.3 Spectral Regularization for Token Alignment
While our method introduces token-level perturbations to simulate distribution shifts, the supervision derived from preference pairs $(y_{w},y_{r})$ is static. This discrepancy between adaptive input representations and fixed supervision may encourage the model to learn distribution-specific artifacts, especially under strong alignment constraints (Fu et al., 2025; Chowdhury et al., 2024).
In practice, semantic alignment does not require strict token-level correspondence. Enforcing fine-grained constraints may reintroduce spurious correlations that our min–max strategy aims to mitigate (Zhou et al., 2024; Tian et al., 2025). To address this, we propose frequency-domain alignment, where local token perturbations translate into smooth variations in spectral space. This approach ensures semantic consistency between perturbed and inputs without rigid token-wise matching.
Specifically, we extract hidden states for $(x,\varphi(q),y_{w})$ and compare them to those of $(x,q,y_{w})$ and $(x,q,y_{r})$ using the Fast Fourier Transform (FFT) (Cooley & Tukey, 1965). Formally, let $z\in\mathbb{R}^{L\times D}$ denote a sequence of hidden states, where $L$ is the token length. We compute the spectral representation via:
$\mathcal{F}(z)=\left|\text{Re}\left[\sum_{l=0}^{L-1}z_{l}\cdot e^{-2\pi ikl/L}\right]\right|_{2},\ \text{for }k=0,\dots,L{-}1,$ (10)
where the FFT is applied along the token axis and $|\cdot|_{2}$ computes the $\ell_{2}$ norm over real-valued frequency magnitudes. The resulting spectral preference loss is defined as:
$\mathcal{L}_{\text{freq}}=-\log\sigma\Big{(}\beta\cdot\Big{[}\,\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{w})\right)}-\log\frac{\mathcal{F}\left(h_{\theta}(x,\varphi(q),y_{w})\right)}{\mathcal{F}\left(h_{\text{ref}}(x,q,y_{r})\right)}\,\Big{]}\Big{)},$ (11)
where $h_{\theta}(\cdot)$ and $h_{\text{ref}}(\cdot)$ are the hidden states of the policy and reference models. This objective follows the logic of DPO but extends alignment to the spectral domain, improving consistency in frequency-aware representations and reducing hallucinations from overfitting to fixed preferences.
@@ 3.4 Minimization Objective in TARS
$\mathcal{L}_{\text{TARS}}(x,q,\varphi(q),y_{w},y_{r})=\;\mathcal{L}_{\text{DPO}}(x,\varphi(q),y_{w},y_{r})+\lambda\cdot\mathcal{L}_{\text{freq}}(x,q,\varphi(q),y_{w},y_{r}),$ (12)
where $\lambda$ is a weighting coefficient that balances preference alignment and spectral consistency. This formulation encourages the model to preserve causal alignment with preference signals, thereby mitigating spurious correlation.
@@ 4.1 Experiment Details
Experiment Setups. We evaluate our method using the multimodal large language model LLaVA-v1.5 (Liu et al., 2023b) at both 7B and 13B parameter scales. All evaluations are conducted using greedy decoding with a temperature of 0 to ensure deterministic outputs and reproducibility.
To enable fair comparison, we align our training configuration with the most data-efficient preference optimization baselines. Specifically, we randomly sample 4.8k instances from the RLHF-V-Dataset (Yu et al., 2024), consistent with OPA-DPO (Yang et al., 2025), and adopt the same training strategy as CHiP-DPO (Fu et al., 2025). All models are trained on eight NVIDIA A100 (80GB) GPUs. We set $\alpha=1$ in Equation 5 and $\beta=1$ in Equation 11 for preference optimization. We implement $\varphi(\cdot)$ using both token masking (Mask) and replacement (Replace) strategies in Equation 8, and set the perturbation constraint strength to $\omega=0.1$ in the adversarial min-max formulation Equation 9. We use a frequency-domain loss weight of $\lambda=0.1$ in Equation 12. Full implementation details and ablation studies are provided in the Appendix.
Evaluation Benchmark. We evaluate TARS across both generative and discriminative hallucination benchmarks to ensure that hallucination mitigation does not come at the cost of factual grounding. Our evaluation framework includes four established benchmarks:
1) AMBER Wang et al. (2023) (Generative): A comprehensive 1004-image benchmark targeting fine-grained evaluation of hallucinations. In line with prior works Yang et al. (2025); Fu et al. (2025), we evaluate only the generative subset using the official codebase. Metrics include CHAIR Rohrbach et al. (2018), object coverage (Cover), response-level hallucination rate (Hal-Rate), and alignment with human cognition (Cog).
2) MMHal (Sun et al., 2023) (Generative): A VQA benchmark with 96 real-world scenarios, evaluated using GPT-4V feedback to measure overall scores and hallucination rates (Hal-Rate).
3) OBJHal (Yu et al., 2024) (Generative): A 300-instance benchmark evaluating hallucinations in captioning. We report hallucination rates at the response level (CRs) and object mention level (CRi).
4) POPE (Li et al., 2023) (Discriminative): A 9000-sample binary VQA benchmark designed to assess object hallucination along the textual axis through yes/no questions.
@@
(1) Advanced multimodal foundation models: Intern-VL2.5-7B (Chen et al., 2024e), Qwen-VL2.5-8B (Bai et al., 2025), DeepSeek-VL2-27B (Wu et al., 2024), and GPT-4o (Hurst et al., 2024) serve as reference benchmarks.
(2) LLaVA-v1.5 with RL techniques: We evaluate multiple RL-based approaches applied to both the 7B and 13B variants of LLaVA-v1.5, including RLHF (Sun et al., 2024), RLAIF (Yu et al., 2025), HALVA (Sarkar et al., 2025a), as well as three state-of-the-art methods based on direct preference optimization (DPO): DPO (Pi et al., 2024), CHiP-DPO (Fu et al., 2025), and OPA-DPO (Yang et al., 2025). A comparison of algorithmic properties is provided in Table 3.
Table 1: Comparison across hallucination benchmarks. We evaluate SOTA MLLMs as reference baselines, denoted by $\S$. For algorithms with available checkpoints, results from re-testing are marked with $\dagger$; for those without, we reproduce results using settings from (Fu et al., 2025; Li et al., 2024), denoted by $\ddagger$. All experiments use greedy decoding with temperature set to 0 for consistency and reproducibility. Bold denotes the best performance, and underlined denotes the second-best.
@@
Intern-VL2.5-7B (Chen et al., 2024e)§ 7.9 54.7 37.1 3.2 3.54 0.26 - - 36.0 9.1
Qwen-VL2.5-8B (Bai et al., 2025)§ 4.6 54.6 21.1 1.3 3.29 0.27 - - 40.7 8.6
DeepSeek-VL2-27B (Wu et al., 2024)§ 2.4 56.6 16.3 0.9 2.84 0.27 - - 10.0 7.0
GPT-4o (Hurst et al., 2024)§ 2.5 60.9 17.6 0.8 3.87 0.24 - - 29.3 6.7
LLaVA-v1.5-7B (Liu et al., 2023b)§ 7.6 51.7 35.4 4.2 2.02 0.61 80.0 61.8 54.0 15.8
+ RLHF (Sun et al., 2024) 8.3 52.2 41.8 4.5 1.93 0.67 82.0 69.3 56.0 15.2
+ RLAIF (Yu et al., 2025) 3.0 50.3 16.5 1.0 2.89 0.42 88.1 88.0 13.7 4.2
+ HALVA (Sarkar et al., 2025a) 6.9 52.8 33.2 3.5 2.12 0.59 87.5 79.6 47.3 14.6
+ DPO (Li et al., 2024) 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ CHiP-DPO (Fu et al., 2025) 2.9 57.3 19.9 1.0 2.32 0.57 81.1 91.8 7.3 4.3
+ OPA-DPO (Yang et al., 2025) 2.7 47.4 12.5 0.9 2.78 0.46 87.4 86.2 13.3 4.5
@@
LLaVA-v1.5-13B (Liu et al., 2023b)§ 6.7 52.1 32.5 3.5 2.39 0.53 74.6 55.2 50.0 14.5
+ RLHF (Sun et al., 2024) 7.1 51.4 36.3 3.6 2.10 0.67 83.6 71.2 46.7 11.6
+ HALVA (Sarkar et al., 2025a) 6.5 53.4 30.1 3.3 2.28 0.56 86.8 75.6 42.7 12.1
+ DPO (Li et al., 2024) 4.1 56.7 24.3 2.2 2.48 0.50 85.2 84.3 19.0 7.2
+ CHiP-DPO (Fu et al., 2025) 3.8 58.6 20.8 1.7 2.70 0.46 86.6 74.9 30.0 6.2
+ OPA-DPO (Yang et al., 2025) 2.8 48.4 13.5 1.0 3.02 0.40 87.2 80.7 18.3 5.1
@@ 4.2 Evaluation on Hallucination Benchmarks
Table 1 presents the performance comparison on four multimodal hallucination benchmarks. For TARS, we employ two perturbation strategies: token masking and synonym replacement (detailed analysis provided in the Appendix). TARS consistently mitigates hallucinations, operates under minimal supervision, scales well, preserves factual accuracy, and remains competitive with proprietary models. Our analysis reveals the following key findings:
@@
(2) Data and supervision efficiency. TARS achieves strong hallucination mitigation without relying on expert-labeled feedback or high-resource teacher models. As shown in Table 3, it uses only 4.8k public preference samples, matching the data budget of OPA-DPO, yet achieves superior performance on AMBER (13B) by improving object coverage from 48.4% to 59.8% (+11.4 pp) and reducing the hallucination rate from 13.5% to 12.5% (–1.0 pp).
(3) Scalability. TARS exhibits consistent performance as model capacity increases. From 7B to 13B, CHAIR improves from 2.4 to 2.1 (-0.3 pp) and hallucination rate drops from 13.2% to 12.5% (-0.7 pp). TARS-13B also surpasses all 13B baselines, confirming the scalability.
(4) Factual consistency. TARS effectively suppresses hallucinations without impairing factual understanding. It achieves 88.7% accuracy on POPE (+8.7 pp over LLaVA-7B) in fine-grained visual reasoning and reduces object-level hallucination on OBJHal to 3.2% (–1.1 pp).
(5) Competitiveness with proprietary models. TARS delivers performance comparable to industrial-scale MLLMs. At 13B, it approaches GPT-4o in AMBER Cover (59.8% vs. 60.9%) and surpasses it in Hal-Rate (12.7% vs. 17.6%), while also outperforming DeepSeek-VL2-27B.
Figure 4: Comparison of average scores across question categories on the MMHal benchmark.
Table 2: Ablation results of token-level perturbation (TP), cross-modal alignment score (CAS), and spectral preference alignment (SPA) under 7B scale.
@@ 4.3 Ablation Analyses on Component
We analyze the contribution of key TARS components through ablations in Figure 5, focusing on three elements:
(1) Token-level perturbation (TP) in Equation 6, which introduces distributional shifts and proves essential for revealing token-level vulnerabilities, its removal increases Cog from 0.4 to 2.5.
(2) Cross-modal alignment score (CAS) in Equation 9, which targets visually agnostic tokens to preserve semantic fidelity—its absence leads to a 4.5-point increase in hallucination and a 0.9 rise in Cog, indicating weaker suppression of spurious correlations.

### v4

@@ 3.3 Spectral Regularization for Token Alignment
Token-level perturbation introduces distribution shifts, yet the supervision from preference pairs $(y_{w},y_{r})$ is static. This discrepancy may cause the model to learn distribution-specific artifacts under strong alignment constraints [CHIP, chowdhuryprovably].
We align in the frequency domain rather than in the spatial domain (e.g., $\ell_{2}$). The key motivation is that a spatial loss $\|z-z^{\prime}\|_{2}$ penalizes every positional deviation equally; when specific tokens are perturbed, this rigid constraint reintroduces spurious positional correlations [zhou2024explore, tianblack]. In contrast, by the shift property of the Discrete Fourier Transform, perturbing a single token $z_{l}$ affects every frequency bin $k$ by an additive term proportional to $e^{-2\pi ikl/L}$, whose magnitude is bounded regardless of $l$. The spectral representation thus absorbs local perturbations into a smooth global envelope, preserving dominant low-frequency semantics while tolerating position-specific noise [cooley1965algorithm]. [^3] A formal theoretical analysis of spectral alignment under token-level adversarial perturbation, including proofs of energy dispersion and bounded semantic shift via low-frequency masking, is provided in the Appendix, Section 9.3.
Concretely, we extract hidden states for $(x,\varphi(q),y_{w})$ and contrast them with $(x,q,y_{w})$ and $(x,q,y_{r})$. Let $z\in\mathbb{R}^{L\times D}$ denote a hidden-state sequence. The spectral representation is:
$\footnotesize\mathcal{F}(z)=\left|\text{Re}\left[\sum_{l=0}^{L-1}z_{l}\cdot e^{-2\pi ikl/L}\right]\right|_{2},\ \text{for }k=0,\dots,L{-}1,$ (10)
where the FFT is applied along the token axis, $\text{Re}[\cdot]$ extracts the real part, and $|\cdot|_{2}$ yields a scalar spectral summary. The spectral preference loss is:
$\mathcal{L}_{\text{freq}}=-\log\sigma\Big(\beta\big[\log\frac{\mathcal{F}(h_{\theta}(x,\varphi(q),y_{w}))}{\mathcal{F}(h_{\text{ref}}(x,q,y_{w}))}-\log\frac{\mathcal{F}(h_{\theta}(x,\varphi(q),y_{r}))}{\mathcal{F}(h_{\text{ref}}(x,q,y_{r}))}\big]\Big).$ (11)
Here $h_{\theta}(\cdot)$ and $h_{\text{ref}}(\cdot)$ are hidden states from the policy and reference models, and $\beta$ is a scaling temperature. This objective extends DPO alignment to the spectral domain, improving frequency-aware consistency and reducing hallucinations from overfitting to fixed preferences.
@@ 3.4 Minimization Objective in TARS
$\footnotesize\mathcal{L}_{\text{TARS}}(x,q,\varphi(q),y_{w},y_{r})=\;\mathcal{L}_{\text{DPO}}(x,\varphi(q),y_{w},y_{r})+\lambda\cdot\mathcal{L}_{\text{freq}}(x,q,\varphi(q),y_{w},y_{r}).$ (12)
where $\lambda$ is a weighting coefficient that balances preference alignment and spectral consistency. This joint formulation explicitly encourages the model to preserve causal alignment with preference signals under adversarial perturbation, thereby mitigating spurious correlation. Extended ablation studies on the sensitivity of $\omega$ and $\lambda$ are presented in Appendix, Section 9.
@@ 4.1 Experiment Details
Experiment Setups. We evaluate our approach on the multimodal LLM LLaVA-v1.5 [llava_origin] at both 7B and 13B scales, and on Muffin-13B [yu2023reformulating, CHIP] (Appendix Section 11). All methods are performed with greedy decoding and a temperature of 0.
For fair comparison, we carefully align our training configuration with the most data-efficient preference optimization baselines. Specifically, we randomly sample 4.8k instances from the RLHF-V-Dataset [yu2024rlhf], consistent with OPA-DPO [OPA_DPO], and adopt the same training strategy as CHiP-DPO [CHIP]. All models are trained on eight NVIDIA A100 (80GB) GPUs with identical hyperparameters. We set $\alpha=1$ in Eq.˜5 and $\beta=1$ in Eq.˜11 for preference optimization. We implement $\varphi(\cdot)$ using both token masking (Mask) and replacement (Replace) strategies in Eq.˜8, and set the perturbation constraint strength to $\omega=0.1$ in the adversarial min-max formulation Eq.˜9. We use a frequency-domain loss weight of $\lambda=0.1$ in Eq.˜12. Full implementation details and ablation studies are reported in the Appendix. [^4] See Appendix, Sections 7 and 9 for configurations and hyperparameter ablations.
Evaluation Benchmarks. We evaluate TARS across four established hallucination benchmarks spanning generative and discriminative settings: AMBER [wang2023amber] for fine-grained generative hallucination (CHAIR [rohrbach2018object], Cover, Hal-Rate, Cog), MMHal [sun2023aligning] for VQA hallucination scored by GPT-4V, OBJHal [yu2024rlhf] for captioning hallucination (response-level CRs and object-level CRi), and POPE [li2023evaluating] for binary discriminative object hallucination. Detailed benchmark descriptions and evaluation protocols are provided in Appendix, Section 7.
@@
(1) Advanced multimodal foundation models: Intern-VL2.5-7B [chen2024expanding], Qwen-VL2.5-8B [bai2025qwen2], DeepSeek-VL2-27B [wu2024deepseek], and GPT-4o [hurst2024gpt].
(2) LLaVA-v1.5 with RL techniques: We evaluate multiple RL-based approaches applied to both the 7B and 13B variants of LLaVA-v1.5, including RLHF [RLHF], RLAIF [RLAIF], HALVA [HALVA], as well as three state-of-the-art methods based on direct preference optimization (DPO): DPO [DPO_MLLMs], CHiP-DPO [CHIP], and OPA-DPO [OPA_DPO]. A comparison of algorithmic properties is provided in Tab.˜4.
Table 1: Comparison across benchmarks. We evaluate SOTA MLLMs as references, denoted by $\S$. For algorithms with available checkpoints, re-tested results are marked with $\dagger$; for those without, we reproduce results using settings from [CHIP, li2024multi], denoted by $\ddagger$. Bold denotes the best performance, and underlined denotes the second-best.
@@
Intern-VL2.5-7B [chen2024expanding]§ 7.9 54.7 37.1 3.2 3.54 0.26 - - 36.0 9.1
Qwen-VL2.5-8B [bai2025qwen2]§ 4.6 54.6 21.1 1.3 3.29 0.27 - - 40.7 8.6
DeepSeek-VL2-27B [wu2024deepseek]§ 2.4 56.6 16.3 0.9 2.84 0.27 - - 10.0 7.0
GPT-4o [hurst2024gpt]§ 2.5 60.9 17.6 0.8 3.87 0.24 - - 29.3 6.7
LLaVA-v1.5-7B [llava_origin]§ 7.6 51.7 35.4 4.2 2.02 0.61 80.0 61.8 54.0 15.8
+ RLHF [RLHF] 8.3 52.2 41.8 4.5 1.93 0.67 82.0 69.3 56.0 15.2
+ RLAIF [RLAIF] 3.0 50.3 16.5 1.0 2.89 0.42 88.1 88.0 13.7 4.2
+ HALVA [HALVA] 6.9 52.8 33.2 3.5 2.12 0.59 87.5 79.6 47.3 14.6
+ DPO [li2024multi] 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ CHiP-DPO [CHIP] 2.9 57.3 19.9 1.0 2.32 0.57 81.1 91.8 7.3 4.3
+ OPA-DPO [OPA_DPO] 2.7 47.4 12.5 0.9 2.78 0.46 87.4 86.2 13.3 4.5
@@
LLaVA-v1.5-13B [llava_origin]§ 6.7 52.1 32.5 3.5 2.39 0.53 74.6 55.2 50.0 14.5
+ RLHF [RLHF] 7.1 51.4 36.3 3.6 2.10 0.67 83.6 71.2 46.7 11.6
+ HALVA [HALVA] 6.5 53.4 30.1 3.3 2.28 0.56 86.8 75.6 42.7 12.1
+ DPO [li2024multi] 4.1 56.7 24.3 2.2 2.48 0.50 85.2 84.3 19.0 7.2
+ CHiP-DPO [CHIP] 3.8 58.6 20.8 1.7 2.70 0.46 86.6 74.9 30.0 6.2
+ OPA-DPO [OPA_DPO] 2.8 48.4 13.5 1.0 3.02 0.40 87.2 80.7 18.3 5.1
@@ 4.2 Evaluation on Hallucination Benchmarks
Tab.˜1 presents results across four hallucination benchmarks. We adopt token masking and synonym replacement as perturbation strategies (extended results on Muffin-13B are in Appendix, Section 11).
@@
(2) Strong data efficiency under limited supervision. Using only 4.8k public preference samples without expert feedback (Tab.˜4), TARS outperforms OPA-DPO on AMBER-13B in both Cover (+11.4 pp) and Hal-Rate (–1.0 pp).
(3) Scalability across model sizes. From 7B to 13B, CHAIR improves from 2.4 to 2.1, Hal-Rate drops from 13.2% to 12.5%, and Cog stays below 0.7. TARS-13B consistently surpasses all 13B baselines by 1.0–1.5 pp in Hal-Rate.
(4) Competitiveness with larger proprietary models. At 13B, TARS approaches GPT-4o in Cover (59.8% vs. 60.9%) while achieving a notably lower Hal-Rate (12.5% vs. 17.6%), and remains competitive with DeepSeek-VL2-27B in both metrics despite its smaller scale. POPE accuracy reaches 88.7% (+8.7 pp over LLaVA-7B), confirming preserved factual grounding (breakdowns in Appendix, Section 10). [^5] TARS also improves fine-grained discriminative detection (Appendix, Section 10.1) and general multimodal understanding (Appendix, Section 10.2).
Figure 4: Comparison of average scores across question categories on MMHal. TARS achieves consistently higher scores, demonstrating stronger visual grounding.
Table 2: Ablation of token-level perturbation (TP), cross-modal alignment score (CAS), and spectral preference alignment (SPA).
@@ 4.3 Ablation Analyses on Component
We analyze the key TARS components through ablations in Tab.˜2, focusing on three elements:
(1) Token-level perturbation (TP) in Eq.˜6, which introduces distributional shifts and proves essential for revealing token-level vulnerabilities and improving robustness. Removing it substantially increases Cog from 0.4 to 2.5, highlighting its critical role in capturing fine-grained alignment uncertainty.
(2) Cross-modal alignment score (CAS) in Eq.˜9, which targets visually agnostic tokens to preserve semantic fidelity. Its absence leads to a 4.5-point increase in hallucination and a 0.9 rise in Cog, indicating weaker suppression of spurious correlations and reduced visual grounding. A direct comparison between adaptive and random/uniform token selection is provided in the Appendix, Section 9.5, confirming the necessity of the proposed mechanism.
@@
4.4 Ablation Analyses on Preference Scale Impact
We examine how preference data scale influences alignment efficiency (Fig.˜5). TARS consistently outperforms DPO across all scales and exhibits sharper gains in the low-data regime: from 0 to 1.8k examples, the 7B and 13B variants reduce hallucination rates by over 15 pp, indicating that TARS captures the core alignment signal early. Beyond 3.6k examples, marginal gains saturate while performance remains stable, highlighting strong data efficiency for practical scenarios where preference annotations are costly or scarce. This behavior is consistent with the dual-space regularization discussed in Sec.˜3.4: by enforcing alignment in both the output probability space and the hidden representation space, TARS extracts richer supervision from each preference pair, reducing the sample complexity needed to reach competitive performance. In contrast, standard DPO requires substantially more data to compensate for its tendency to overfit to surface-level patterns, as reflected by the persistent hallucination rate gap across all data scales in Fig.˜5.
4.5 Stability of Semantic Representations
We analyze how preference optimization reshapes hidden-state distributions in Fig.˜6. TARS yields a more structured latent space in which hallucinated and preference-aligned representations are clearly separated, whereas DPO produces entangled clusters that interleave hallucinated and preference features, indicating overfitting to superficial signals. Crucially, TARS selectively aligns non-hallucinated responses with preference features while isolating hallucinated content, creating a semantically faithful representation space that reinforces only factually grounded outputs.
The marginal distributions (top and right panels in Fig.˜6) further reveal the mechanism behind this separation. Under DPO, hallucinated and preference-aligned marginals overlap heavily, reflecting a representation space where surface-level patterns dominate over grounded semantics. In contrast, TARS produces near-disjoint marginals for hallucinated content while maintaining tight overlap between non-hallucinated and preference-aligned distributions. This selective clustering indicates that spectral regularization combined with adversarial perturbation shapes a representation geometry that naturally distinguishes factual from hallucinated outputs, without requiring explicit labels during training.
4.6 Training Dynamics Analysis
To validate the min–max formulation (Eq.˜6), we compare training dynamics of standard DPO, random perturbation (RandPert), and TARS (Min-Max) on LLaVA-v1.5-7B over 20 epochs (Fig.˜7; all losses normalized to $[0,1]$). DPO converges to the lowest training loss ($\sim$0.2) but exhibits the worst robust-sample test loss (0.65–0.75), confirming severe overfitting to static preference patterns [DPO_MLLMs]. RandPert shows intermediate test performance but with heavy oscillation, as indiscriminate perturbation yields inconsistent gradients. TARS maintains a moderately higher training loss (0.35–0.4) yet achieves the lowest test loss on both clean (<<0.3) and robust (<<0.5) splits with stable convergence. This confirms that the min–max formulation acts as an implicit regularizer: adversarial token-level shifts during training prevent premature convergence and yield alignment features that generalize to unseen and perturbed data.
Figure 7: Training dynamics of DPO, RandPert, and TARS (Min-Max); all losses are normalized to $[0,1]$. (a) Training objective: DPO converges fastest but overfits; Min-Max maintains a higher loss reflecting adversarial difficulty. (b) Test performance on clean and robust (perturbed) samples: Min-Max achieves the lowest test loss on both splits, while DPO exhibits a large clean–robust gap indicating poor robustness.
4.7 Comparison with Data Augmentation
[…]
Discussion. Across all experiments, TARS’s improvements manifest systematically across generative (AMBER, OBJHal), discriminative (POPE), and open-ended (MMHal) evaluations. The ablation studies (Tab.˜2) confirm that each component contributes independently, the scale analysis (Fig.˜5) demonstrates strong data efficiency, and the training dynamics (Fig.˜7) validate the regularization effect of the min-max formulation. This convergence of evidence across complementary evaluation axes supports the claim that token-adaptive perturbation with spectral regularization addresses a fundamental limitation of static preference optimization rather than exploiting benchmark-specific biases.
Table 3: Comparison of TARS against data augmentation baselines on four hallucination benchmarks. Paraphrasing denotes back-translation-based response diversification; LLM Aug. denotes LLM-generated preferred responses, where $1\times$ and $5\times$ indicate the expansion factor relative to the original 4.8k preference set. All methods build on standard DPO. TARS is complementary to augmentation strategies.
Method AMBER MMHal POPE OBJHal
[…]
LLaVA-7B-DPO 4.9 56.6 26.4 2.5 2.19 0.61 87.8 82.0 14.0 5.0
+ Paraphrasing 4.5 57.9 24.8 2.3 2.24 0.58 87.9 83.5 13.6 4.8
[…]
+ TARS 2.4 59.6 13.2 0.4 2.48 0.45 88.7 97.5 12.0 3.2
+ TARS & Para. 2.4 60.1 12.8 0.4 2.50 0.45 88.9 97.5 11.6 3.0
[…]
5 Related Work